40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 26 17:24:24 2017, Steve, Update, General, power outage powerOutage.png
    Reply  Thu Dec 28 17:22:02 2017, gautam, Update, General, power outage - CDS recovery 28.png
       Reply  Thu Jan 4 14:39:50 2018, gautam, Update, General, power outage - timing error 
          Reply  Fri Jan 5 21:54:28 2018, rana, Update, General, power outage - timing error huh.png
             Reply  Fri Jan 5 22:19:53 2018, gautam, Update, General, power outage - timing error 
                Reply  Sat Jan 6 18:27:37 2018, gautam, Update, General, power outage - IFO recovery EX_QPD_Quad1_Faulty.pdfTrillium_faulty.pdf
                   Reply  Mon Apr 16 22:09:53 2018, Kevin, Update, General, power outage - BLRM recovery AA.jpgADC.jpg
                      Reply  Tue Apr 17 09:57:09 2018, Steve, Update, General, seismometer interfaces BS_Tril_Intrf-1X5.jpgGurs_Intf-1X1.jpg
                         Reply  Wed Apr 18 20:33:19 2018, Kevin, Update, General, seismometer interfaces front.jpgback.jpgconnectors.jpg
                            Reply  Fri Apr 20 23:36:28 2018, Kevin, Update, PEM, Seismometer BLRMs front.jpgback.jpgdetail.jpg
                               Reply  Tue Apr 24 21:19:08 2018, Kevin, Update, PEM, Seismometer BLRMs modified_schematic.pdftop.jpgbottom.jpgBS_Seis_PSD.pdf
                                  Reply  Wed Apr 25 17:44:39 2018, Arnold, Update, PEM, PEM Anti-Alias wiring 
                                     Reply  Thu Apr 26 09:35:49 2018, Kevin, Update, PEM, PEM Anti-Alias wiring board.jpgSeismometerPSD.pdf
                                        Reply  Thu Apr 26 20:22:21 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections setup.pdfseismometerASD.pdf
                                           Reply  Wed May 16 21:02:22 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections ADC_noise.pdf
                                        Reply  Thu Jun 14 15:24:32 2018, Steve, Update, PEM, ADC DAC In Line Test Boards are in ADC_DAC_in_(1).JPG
    Reply  Tue Jan 2 15:43:35 2018, Steve, Update, VAC, pumpdown after power outage pumpdown_from_44_mTorr.png
       Reply  Wed Feb 7 15:50:42 2018, Steve, Update, VAC, IFO pressure monitoring CC_analog.png
Message ID: 13493     Entry time: Thu Dec 28 17:22:02 2017     In reply to: 13492     Reply to this: 13503
Author: gautam 
Type: Update 
Category: General 
Subject: power outage - CDS recovery 
  1. I had to manually reboot c1lsc, c1sus and c1ioo.
  2. I edited the line in /etc/rt.sh (specifically, on FB /diskless/root.jessie/etc/rt.sh) that lists models running on a given frontend, to exclude c1dnn and c1oaf, as these are the models that have been giving us most trouble on startup. After this, I was able to bring back all models on these three machines using rtcds restart --all. The original line in this file has just been commented out, and can be restored whenever we wish to do so.
  3. mx_stream processes are showing failed status on all the frontends. As a result, the daqd processes are still not working. Usual debugging methods didn't work.
  4. Restored all sus dampings.
  5. Slow computers all seem to be responsive, so no action was required there.
  6. Burtrestored c1psl to solve the "sticky slider" problem, relocked PMC. I didn't do anything further on the PSL table w.r.t. the manual beam block Steve has placed there till the vacuum situation returns to normal.

@Steve: I noticed that we are down to our final bottle of N2, not sure if it will last till 2 Jan which is presumably when the next delivery will come in. Since V1 is closed and the PSL beam is blocked, perhaps this doesn't matter.

from Steve: there are spare full N2 bottles at the south end outside and inside. I replaced the N2 on Sunday night. So the system should be Ok as is.

I also hard-rebooted megatron and optimus as these were unresponsive to ping.

*Seems like the mx_stream errors were due to the mx process not being started on FB. I could fix this by running sudo systemctl start mx on FB. After which I ran sudo systemctl restart daqd_*. But the DC errors persist - not sure how to fix this. Elogging suggests that "0x4000" errors are connected to timing problems on FB, but restarting the ntp service on FB (which is the suggested fix in said elogs) didn't fix it. Also unsure if mx process is supposed to automatically start on FB at startup.

Attachment 1: 28.png  77 kB  | Hide | Hide all
28.png
ELOG V3.1.3-